Find three consecutive odd integers such that the largest decreased by three times the second is the same as the difference of the smallest and 47.
Answer by checkley77(7072)

Let x, x+2 & x+4 be the 3 odd consecutive intergers.
(x+4)-3(x+2)=x-47
x+4-3x-6=x-47
x-3x-x=-47-4+6
-3x=-45
x=-45/-3
x=15 ans. for the smallest integer.
15+2=17 ans. for the middle integer.
15+4=19 ans. for the largest integer.
Proof:
19-3*17=15-47
19-51=-32
-32=-32